OpenStudy (anonymous):

Find the quotient. Show all steps please. 25a^2 - 16b^2 / 5a - 4b

myininaya (myininaya):

trying factoring the top

myininaya (myininaya):

then something will cancel

myininaya (myininaya):

and then you will be done

jimthompson5910 (jim_thompson5910):

\[\large \frac{25a^2-16b^2}{5a-4b}\] \[\large \frac{(5a-4b)(5a+4b)}{5a-4b}\] \[\large \frac{\cancel{(5a-4b)}(5a+4b)}{\cancel{5a-4b}}\] \[\large 5a+4b\] So \[\large \frac{25a^2-16b^2}{5a-4b}=5a+4b\]

myininaya (myininaya):

like that

OpenStudy (anonymous):

it looks great :)